Sixteen liters of an ideal gas at a temperature of 500 kelvins has a pressure of 100 newtons per square meter. If the volume were increased to 20 liters and the temperature reduced to 400 kelvins, what would be the pressure?
Answer by stanbon(57940)

Sixteen liters of an ideal gas at a temperature of 500 kelvins has a pressure of 100 newtons per square meter.
---
Volume = k[t/p]
---
Find "k":
16 = k[500/100]
---
k = 16/5
----
Equation:
V = (16/5)[t/p]
---
If the volume were increased to 20 liters and the temperature reduced to 400 kelvins, what would be the pressure?
20 = (16/5)[400/p]
pressure = (16/5)[500/20]
---
Pressure = 16 neutons per sq. meter
==============
Cheers,
Stan H.
==============